# On an inequality of Lin, Kim and Hsieh and Strong Subadditivity

### v1

@@ 1 Introduction
Let $\mathcal{H}_{1},\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ be finite dimensional Hilbert spaces. For a density matrix $\rho_{123}$ on $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}$, let $\rho_{12}:=\textnormal{Tr}_{3}[\rho_{123}]$, the partial trace of $\rho_{123}$ over $\mathcal{H}_{3}$, and $\rho_{1}:=\textnormal{Tr}_{2}[\rho_{12}]=\textnormal{Tr}_{23}[\rho_{123}]$, etc. In 1968, Lanford and Robinson [6] conjectured that the von Neumann entropies of $\rho_{123}$, $\rho_{2}$, $\rho_{12}$ and $\rho_{23}$ we related by
@@
In 1970, while the conjecture was still open, Araki and Lieb [3] showed by an elementary purification argument that (1) hold for all tripartite density matrices $\rho_{123}$ if and only if
@@
holds for all tripartite density matrices $\rho_{123}$. They did prove a weaker form of (1), with $S(\rho_{2})$ replaced by $log(\textnormal{Tr}[\rho_{2}^{2}])$, and showed that this sufficed to prove the thermodynamic limit that Lanford and Robinson had sought to prove. However, (1) is of interest for many reasons, and work towards proving it continued.
@@ 2 Proof of the inequality of Lin, Kim and Hsieh
Now taking $\epsilon$ to zero, we recover (8) for pure states $\sigma_{23}$ such that $\sigma_{3}$ is invertible. Such pure states are dense by (3) and (4), and again, since any density matrix on $\mathcal{H}_{2}\otimes\mathcal{H}_{3}$ is a convex combination of pure states, it follows that (8) is valid for all density matrices $\sigma_{12}$. ∎
@@
Let $\{\Phi_{j}\}_{1\leqslant j\leqslant d_{3}d_{3}}$ be an orthonormal basis for $\mathcal{H}_{3}\otimes\mathcal{H}_{3}$ in which $\Phi_{1}=\Phi$. The general unit vector $\varphi\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}$ has the form $\varphi=\sum_{\ell=1}^{d_{2}d_{3}}{\bf w}_{\ell}\otimes\Phi_{\ell}$ where $\sum_{\ell=1}^{d_{2}d_{3}}\|{\bf w}_{\ell}\|^{2}=1$. Then with $B:=\rho_{1}\otimes X_{23}(\epsilon)^{-1}$,
@@
We must compare this with $\langle\varphi,A\varphi\rangle$, where $A:=\rho_{12}\otimes\sigma_{3}^{-1}$. Recall that for all $a,b,\delta>0$, $ab\leqslant\frac{1}{2}(\delta a^{2}+\delta^{-1}b)$. Hence for any $\delta>0$,
@@
$\langle{\bf w}_{1}\otimes\Phi,A{\bf w}_{1}\otimes\Phi\rangle=\langle{\bf w}_{1},\rho_{1}{\bf w}_{1}\rangle\ .$ (16)
Once this identity is established, choose $\delta=\sqrt{\epsilon}/d_{2}d_{3}$ and then (15) becomes
@@
It remains to prove (16). Let ${\displaystyle\Psi=\sum_{j=1}^{d_{1}}\lambda_{j}^{1/2}{\bf u}_{j}\otimes{\bf v}_{j}}$ and ${\displaystyle\Phi=\sum_{j=1}^{d_{3}}\mu_{k}^{1/2}{\bf x}_{k}\otimes{\bf y}_{k}}$ be Schmidt decompositions of $\Psi$ and $\Phi$. Then $\rho_{1}=\sum_{j=1}^{d_{1}}\lambda_{j}|{\bf u}_{j}\rangle\langle{\bf u}_{j}|$ , $\sigma_{3}=\sum_{j=1}^{d_{3}}\mu_{k}|{\bf y}_{k}\rangle\langle{\bf y}_{k}|$ and
$A=\sum_{j,j^{\prime},k}\lambda_{j}^{1/2}\lambda_{j^{\prime}}^{1/2}\mu_{k}^{-1}|{\bf u}_{j}\otimes{\bf v}_{j}\otimes{\bf y}_{k}\rangle\langle{\bf u}_{j^{\prime}}\otimes{\bf v}_{j^{\prime}}\otimes{\bf y}_{k}|\ .$ (18)
Writing out $\langle{\bf w}_{1}\otimes\Phi,A{\bf w}_{1}\otimes\Phi\rangle$ using the Schmidt decomposition for $\Phi$ yields the sum
@@
$\sum_{j,j^{\prime},k}\lambda_{j}^{1/2}\lambda_{j^{\prime}}^{1/2}\langle{\bf w}_{1},{\bf u}_{j}\rangle\langle{\bf u}_{j^{\prime}},{\bf w}_{1}\rangle\langle{\bf v}_{j^{\prime}},{\bf x}_{k}\rangle\langle{\bf x}_{k},{\bf v}_{j}\rangle=\langle{\bf w}_{1},\rho_{1}{\bf w}_{1}\rangle\ .$
The original proof of strong subadditivity [6] showed that strong subadditivity, the joint convexity of the relative entropy, and the concavity of the conditional entropy were all equivalent, and then used the convexity results from [5] to prove strong subadditivity. It is remarkable that the proof of strong subadditivity using Theorem 1.1 sidesteps deep convexity issues altogether, and indeed, using the equivalence results, provides simple proofs of the joint convexity of the relative entropy, and the concavity of the conditional entropy.

### v2

@@ 1 Introduction
Let $\mathcal{H}_{1},\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ be finite dimensional Hilbert spaces. For a density matrix $\rho_{123}$ on $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}$, let $\rho_{12}:=\textnormal{Tr}_{3}[\rho_{123}]$, the partial trace of $\rho_{123}$ over $\mathcal{H}_{3}$, and $\rho_{1}:=\textnormal{Tr}_{2}[\rho_{12}]=\textnormal{Tr}_{23}[\rho_{123}]$, etc. In 1968, Lanford and Robinson [6] conjectured that the von Neumann entropies of $\rho_{123}$, $\rho_{2}$, $\rho_{12}$ and $\rho_{23}$ are related by
@@
In 1970, while the conjecture was still open, Araki and Lieb [3] showed by an elementary purification argument that (1) is valid for all tripartite density matrices $\rho_{123}$ if and only if
@@
holds for all tripartite density matrices $\rho_{123}$. They did prove a weaker form of (1), with $S(\rho_{2})$ replaced by $\log(\textnormal{Tr}[\rho_{2}^{2}])$, and showed that this sufficed to prove the thermodynamic limit that Lanford and Robinson had sought to prove. However, (1) is of interest for many reasons, and work towards proving it continued.
@@ 2 Proof of the inequality of Lin, Kim and Hsieh
Now taking $\epsilon$ to zero, we recover (8) for pure states $\sigma_{23}$ such that $\sigma_{3}$ is invertible. Such pure states are dense by (3) and (4), and again, since any density matrix on $\mathcal{H}_{2}\otimes\mathcal{H}_{3}$ is a convex combination of pure states, it follows that (8) is valid for all density matrices $\sigma_{23}$. ∎
@@
Let $\{\Phi_{j}\}_{1\leqslant j\leqslant d_{2}d_{3}}$ be an orthonormal basis for $\mathcal{H}_{2}\otimes\mathcal{H}_{3}$ in which $\Phi_{1}=\Phi$. The general unit vector $\varphi\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}$ has the form $\varphi=\sum_{\ell=1}^{d_{2}d_{3}}{\bf w}_{\ell}\otimes\Phi_{\ell}$ where $\sum_{\ell=1}^{d_{2}d_{3}}\|{\bf w}_{\ell}\|^{2}=1$. Then with $B:=\rho_{1}\otimes X_{23}(\epsilon)^{-1}$,
@@
We must compare this with $\langle\varphi,A\varphi\rangle$, where $A:=\rho_{12}\otimes\sigma_{3}^{-1}$. Recall that for all $a,b,\delta>0$, $ab\leqslant\frac{1}{2}(\delta a^{2}+\delta^{-1}b^{2})$. Hence for any $\delta>0$,
@@
$\langle{\bf w}_{1}\otimes\Phi,A{\bf w}_{1}\otimes\Phi\rangle\leqslant\langle{\bf w}_{1},\rho_{1}{\bf w}_{1}\rangle\ .$ (16)
Accepting this, choose $\delta=\sqrt{\epsilon}/d_{2}d_{3}$ and then (15) becomes
@@
To prove (16), let ${\displaystyle\Psi=\sum_{j=1}^{d_{1}}\lambda_{j}^{1/2}{\bf u}_{j}\otimes{\bf v}_{j}}$ and ${\displaystyle\Phi=\sum_{k=1}^{d_{3}}\mu_{k}^{1/2}{\bf x}_{k}\otimes{\bf y}_{k}}$ be Schmidt decompositions of $\Psi$ and $\Phi$. Then $\rho_{1}=\sum_{j=1}^{d_{1}}\lambda_{j}|{\bf u}_{j}\rangle\langle{\bf u}_{j}|$ , and ${\displaystyle A=\sum_{j,j^{\prime},k}\lambda_{j}^{1/2}\lambda_{j^{\prime}}^{1/2}\mu_{k}^{-1}|{\bf u}_{j}\otimes{\bf v}_{j}\otimes{\bf y}_{k}\rangle\langle{\bf u}_{j^{\prime}}\otimes{\bf v}_{j^{\prime}}\otimes{\bf y}_{k}|}$. Writing out $\langle{\bf w}_{1}\otimes\Phi,A{\bf w}_{1}\otimes\Phi\rangle$ using the Schmidt decomposition for $\Phi$ yields the sum
@@
$\sum_{j,j^{\prime},k}\lambda_{j}^{1/2}\lambda_{j^{\prime}}^{1/2}\langle{\bf w}_{1},{\bf u}_{j}\rangle\langle{\bf u}_{j^{\prime}},{\bf w}_{1}\rangle\langle{\bf v}_{j^{\prime}},{\bf x}_{k}\rangle\langle{\bf x}_{k},{\bf v}_{j}\rangle\ .$
Define the $d_{1}\times d_{1}$ matrix $M$ with entries $M_{j^{\prime},j}:=\sum_{k=1}^{d_{3}}\langle{\bf v}_{j^{\prime}},{\bf x}_{k}\rangle\langle{\bf x}_{k},{\bf v}_{j}\rangle$ and the vector ${\bf a}\in\mathbbm{C}^{d_{1}}$ with entries $a_{j}:=\lambda_{j}^{1/2}\langle{\bf w}_{1},{\bf u}_{j}\rangle$. Then $0<M\leqslant\mathds{1}$ (with equality on the right if $d_{3}=d_{2}$), and therefore, $\langle{\bf w}_{1}\otimes\Phi,A{\bf w}_{1}\otimes\Phi\rangle=\langle{\bf a},M{\bf a}\rangle\leqslant\|{\bf a}\|^{2}=\langle{\bf w}_{1},\rho_{1}{\bf w}_{1}\rangle$. ∎
The original proof of strong subadditivity [6] showed that strong subadditivity, the joint convexity of the relative entropy, and the concavity of the conditional entropy were all equivalent, and then used the convexity results from [5] to prove strong subadditivity. The proof of strong subadditivity using Theorem 1.1 sidesteps deep convexity issues altogether, and indeed, using the equivalence results, provides simple proofs of the joint convexity of the relative entropy, and the concavity of the conditional entropy. We thank Trung Nghia Nguyen for pointing out a flaw in an earlier proof of the main lemma.
We close by showing that there are no cases of equality for the operator inequality (8).
Theorem 2.2.
If ${\rm dim}(\mathcal{H}_{2})>1$, there are no density matrices $\rho_{12}$ and $\sigma_{23}$ for which there is equality in (8).
Proof.
Suppose there is equality in (8). Discarding the kernel of $\sigma_{3}$ from $\mathcal{H}_{3}$ as needed, we may suppose $\sigma_{3}$ is invertible. (Note that as operators on $\mathcal{H}_{2}\otimes\mathcal{H}_{3}$, ${\rm ker}(\sigma_{3})\subseteq{\rm ker}(\sigma_{23})$). Then with equality in (8), $\sigma_{23}$ is in fact invertible, and of course $\rho_{1}$ is invertible whenever $\rho_{12}$ is invertible.
Define $X=\sigma_{3}^{-1/2}\sigma_{23}\sigma_{3}^{-1/2}$ and $Y=\rho_{1}^{-1/2}\rho_{12}\rho^{-1/2}$. Then as operators on $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}$,
$X=Y^{-1}$ (18)
Since $X$ acts trivially on $\mathcal{H}_{1}$ and $Y$ acts trivially on $\mathcal{H}_{3}$, there are $A,B\in\mathcal{B}^{+}(\mathcal{H}_{2})$ such that $X=\mathds{1}\otimes A\otimes\mathds{1}$ and $Y=\mathds{1}\otimes B\otimes\mathds{1}$. From the definitions of $X$ and $Y$, $\textnormal{Tr}_{2}[X]=\textnormal{Tr}_{2}[Y]=\mathds{1}$ on $\mathcal{H}_{1}\otimes\mathcal{H}_{3}$, and this means that $\textnormal{Tr}_{2}[A]=\textnormal{Tr}_{2}[B]=1$. However then $\textnormal{Tr}_{2}[B^{-1}]\geqslant(\textnormal{Tr}_{2}[B])^{-1}$, and equality is only possible if $\mathcal{H}_{2}$ is one dimensional. Hence equality in (8) is not possible when $\mathcal{H}_{2}$ is non-trivial. ∎
3 Declarations
Conflicts of Interest: There is no conflict of interest. Data Availability: there is no data associated to this paper.